# Extracting Interpretable Logic Rules from Graph Neural Networks

### v1

@@ 5 Related Work
The explainability of GNNs remains a relatively underexplored area compared to other neural networks, such as convolutional neural networks (CNNs). Most existing methods focus on providing local input attribution explanations (Pope et al., 2019; Ying et al., 2019; Luo et al., 2020b; Vu & Thai, 2020; Lucic et al., 2022; Tan et al., 2022), similar to attribute-based approaches like Grad-CAM (Selvaraju et al., 2017) used for explaining CNNs. Another line of research focuses on global explanations that capture the overall behavior of models, which is where our approach, LogicXGNN, also belongs. Global explanation methods can be broadly divided into generation-based and concept-based approaches. Generation-based methods, such as XGNN (Yuan et al., 2020), use reinforcement learning agents to generate graph instances that maximize specific model predictions. Similarly, GNNInterpreter (Wang & Shen, 2023) learns a probabilistic generative graph distribution to produce graph patterns that serve as explanations. These approaches are black-boxed and provide only a limited number of explanation patterns for certain classes, whereas LogicXGNN is transparent and can generate any desired instance in a controlled manner, making it highly applicable to fields like drug design. On the other hand, concept-based approaches aim to provide more human-readable and precise explanations by leveraging logical formulas and interpretable concepts. For example, GCneuron (Xuanyuan et al., 2023) identifies predefined concepts, formulated as logical combinations of node degrees and neighborhood properties, associated with specific neurons. Similarly, GLGExplainer (Azzolin et al., 2023) builds on local explanations from PGExplainer (Luo et al., 2020a), maps them to learned concepts and derives logic formulas from these concepts to serve as class explanations. In contrast, LogicXGNN offers a simpler and more intuitive approach, relying solely on decision tree computations rather than complex local explanation methods. It is data-driven, eliminating the need for predefined concepts. Most importantly, while these concept-based approaches typically generate descriptive rules for individual classes without distinguishing between them, LogicXGNN excels in creating discriminative rules that differentiate classes and matches the original model’s accuracy.
6 Conclusion and Future Work
In this work, we present LogicXGNN, a novel framework for extracting interpretable logic from GNNs. LogicXGNN is architecture-agnostic and focuses on extracting global rules rather than instance-level explanations. Unlike existing global methods, it generates discriminative rules that effectively differentiate between classes while achieving accuracy comparable to the original model. Additionally, LogicXGNN can function as a generative model for creating graph instances, demonstrating considerable potential in applications such as drug design. For future work, we identify two key directions to extend LogicXGNN: 1) Enhancing Rule-based Models with Program Synthesis: Since LogicXGNN translates GNN decision-making into logic/program space, we aim to leverage program synthesis techniques and human prior knowledge to further improve the performance of our rule models; 2) Interpretable Approaches for Regression Tasks: Current methodologies of LogicXGNN are tailored for classification tasks. Extending the framework to handle regression problems, such as predicting molecular properties and 3D structures, remains an exciting challenge for future exploration.
Impact Statement
The interpretability of Graph Neural Networks (GNNs) is an essential aspect of their application, particularly in fields like drug discovery, healthcare, and safety-critical systems, where understanding the rationale behind model predictions can have significant consequences. LOGICXGNN tackles the challenge of interpretability by extracting human-readable logic rules from GNNs without relying on predefined concepts, offering a data-driven approach that makes the decision-making process more transparent. This not only enhances our understanding of GNN predictions but also facilitates the discovery of valuable knowledge that may otherwise be overlooked, as demonstrated by experiments on real-world datasets such as MUTAG and BBBP.
One of the key advantages of LOGICXGNN is its potential to replace GNNs as a classifier in scenarios where high fidelity and interpretability are paramount. This is particularly important in settings such as drug design, where model transparency can help prevent harmful or unintended outcomes. Additionally, by enabling the generation of new graph instances in a controlled and transparent manner, LOGICXGNN opens up new possibilities for innovative and reliable applications across various domains.
From an ethical standpoint, LOGICXGNN promotes greater accountability and trust in AI-driven decisions by providing understandable justifications for its predictions. However, it is crucial to acknowledge the potential risks of over-relying on automated systems, particularly in high-stakes contexts. As such, while LOGICXGNN can complement human expertise, it should not replace human oversight, especially when it comes to critical decision-making processes.
Appendix A LogicXGNN as Generative Model
We introduce Algorithm 1 for generating graph instances with LogicXGNN and illustrate the process with a step-by-step real-world example in Figure 5 for clarity.

### v2

@@ 5 Related Work
The explainability of GNNs remains a relatively underexplored area compared to other neural networks, such as convolutional neural networks (CNNs). Most existing methods focus on providing local input attribution explanations [18, 31, 15, 24, 13, 22], similar to attribute-based approaches like Grad-CAM [20] used for explaining CNNs. Another line of research focuses on global explanations that capture the overall behavior of models, which is where our approach, LogicXGNN, also belongs. Global explanation methods can be broadly divided into generation-based and concept-based approaches. Generation-based methods, such as XGNN [32], use reinforcement learning agents to generate graph instances that maximize specific model predictions. Similarly, GNNInterpreter [26] learns a probabilistic generative graph distribution to produce graph patterns that serve as explanations. These approaches are black-boxed and provide only a limited number of explanation patterns for certain classes, whereas LogicXGNN is transparent and can generate any desired instance in a controlled manner, making it highly applicable to fields like drug design. On the other hand, concept-based approaches aim to provide more human-readable and precise explanations by leveraging logical formulas and interpretable concepts. For example, GCneuron [30] identifies predefined concepts, formulated as logical combinations of node degrees and neighborhood properties, associated with specific neurons. Similarly, GLGExplainer [3] builds on local explanations from PGExplainer [14], maps them to learned concepts and derives logic formulas from these concepts. GraphTrail uses symbolic regression to fit a surrogate rule model on subgraph-level concepts as its explanation. In contrast, LogicXGNN offers a simpler and more intuitive approach, relying solely on decision tree computations rather than complex local explanation methods. It is data-driven, eliminating the need for predefined concepts. Most importantly, while these concept-based approaches typically generate descriptive rules for individual classes without distinguishing between them, LogicXGNN excels in creating discriminative rules that differentiate classes and matches the original model’s accuracy.
6 Conclusion, Limitations and Future Work
In this work, we present LogicXGNN, a novel framework for extracting interpretable logic rules from GNNs. LogicXGNN is model-agnostic, efficient, and data-driven. More importantly, it can function as a rule-based classifier and even outperform the original neural models. Its interpretability facilitates knowledge discovery, as demonstrated by its ability to extract detailed and chemically accurate insights that are often overlooked by existing methods. A key advantage of LogicXGNN is its capacity to generate new graph instances in a controlled and transparent manner, offering significant potential for applications such as drug design. However, this also introduces potential risks that must be carefully considered—for instance, the risk of overtrusting the extracted knowledge or generated molcuer graphs without further validation. These risks highlight the need for cautious deployment and expert oversight. While LogicXGNN outperforms state-of-the-art global explanation methods, a performance gap remains compared to the original GNNs on datasets with more complex node interactions and weaker structural regularities, such as IMDB-BINARY. Addressing this limitation is an important direction for future work. We are also interested in exploring program synthesis on top of LogicXGNN to further enhance the performance of the rule-based classifier.
Appendix A Appendix
A.1 Comparison of explanation methods by scope and supported functionalities
Global explanation methods, particularly rule-based ones, are recognized for providing greater explanatory power than local feature attribution methods [34], such as GNNExplainer. For instance, while GNNExplainer computes importance scores for specific substructures such as NO2 groups and carbon rings in a molecule, our rule-based method confirms their essential and stable role across diverse molecules through logical rules. Consistent with prior work on global explanations (e.g., GraphTrail, GLGExplainer), we limit comparisons to global baselines, excluding local methods such as GNNExplainer due to their differing scope.
In summary, we compare our approach with several well-known explanation methods in terms of functionality, as outlined below.
Table 2: Comparison of explanation methods by scope and supported functionalities.
Method Scope Type Classification Knowledge Extraction Graph Generation
[…]
A.2 Dataset details
Mutagenicity [6], NCI1 [25], and BBBP [27] are molecular graph datasets, where nodes represent atoms and edges correspond to chemical bonds. In NCI1, each molecular graph is labeled based on its anticancer activity. The Mutagenicity (MUTAG) dataset contains compounds labeled according to their mutagenic effect on the Gram-negative bacterium Salmonella typhimurium (Label 0 is mutagenic). In the BBBP dataset, molecules are labeled based on their ability to penetrate the blood-brain barrier. BAMultiShapes [31] is a synthetic dataset consisting of 1,000 Barabási-Albert graphs with randomly placed network motifs, including house, grid, and wheel structures. Class 0 contains plain BA graphs and those augmented with one or more motifs, while Class 1 includes graphs enriched with two motif combinations. IMDB-BINARY [16] is a social network dataset where each graph represents a movie and nodes correspond to actors, with edges indicating co-appearances in scenes. We summarize the statistics of the datasets in Table 3.
Table 3: Statistics of the datasets.
BAMultiShapes Mutagenicity BBBP NCI1 IMDB-BINARY
#Graphs 1,000 4,337 2,050 4,110 1,000
Avg. $|\mathcal{V}|$ 40 30.32 23.9 29.87 19.8
Avg. $|\mathcal{E}|$ 87.00 30.77 51.6 32.30 193.1
#Node features 10 14 9 37 0
A.3 Experimental setup
All experiments are conducted on an Ubuntu 22.04 LTS machine equipped with 128 GB of RAM and an AMD EPYC™ 7532 processor. Each dataset is split into training and testing sets using an 80/20 ratio. All experiments are repeated using three different random seeds to ensure robustness.
To demonstrate the model-agnostic nature of LogicXGNN, we employ different GNN architectures across datasets: a 3-layer GCN [10] for BAMultiShapes, a 2-layer GCN for BBBP, a 2-layer GraphSAGE [8] for Mutagenicity, a 3-layer GIN [29] for NCI1, and a 2-layer GAT [23] for IMDB-BINARY.
For training, we use the Adam optimizer with a learning rate of 0.005. Training proceeds for up to 500 epochs, with early stopping after a 100-epoch warm-up if validation accuracy does not improve for 50 consecutive epochs. All explainers are evaluated on the test split.
Additionally, we use the CART algorithm to construct the decision trees in LogicXGNN. For baseline approaches, we adopt the authors’ original implementations and follow their recommended hyperparameters to ensure a fair comparison.
Reproducibility analysis of GLGExplainer
We reproduced the results of GLGExplainer [3] using the official GitHub repository, but encountered several challenges that raise concerns about the fairness of direct comparisons, many of which were also reported by GraphTrail [2]
First, although GLGExplainer relies on PGExplainer [14] to generate local explanations (i.e., important subgraphs), the public codebase only includes precomputed outputs rather than the implementation for invoking PGExplainer. This omission prevents the method from being directly applied—using the authors’ original parameters—to additional datasets beyond those used in the original paper. As a result, we conducted our own trials on other datasets.
Second, GLGExplainer requires prior domain knowledge to learn concept representations for each cluster of local explanations produced by PGExplainer. It also performs dataset-level filtering to exclude graphs that are less relevant to this domain knowledge. For example, in the MUTAG dataset, it relies on the assumption that nitro groups (NO2) and amines (NH2) are indicative of mutagenicity. Based on this assumption, it filters out molecules that do not contain these functional groups, as illustrated in Figure 4.
Figure 4: The screenshot of GLGexplainer github repository.The comment in this code shows that the GLGexplainer remove the graph contains $NH_{2}$
In this code snippet, GLGExplainer excludes the following types of graphs:
  • Graphs whose important subgraphs are nearly as large as the original graph,
  • Graphs with overly simplistic explanations (e.g., single-edge motifs),
  • Graphs containing unique patterns (e.g., the only graph with an NH2 group).
These filtering criteria introduce bias by favoring instances with clear and frequent motifs, potentially overstating the model’s performance. In contrast, our method and GraphTrail make no such assumptions and operate without relying on prior domain knowledge.
Third, while the paper claims to use the elbow method to threshold edge importance scores, we found that for the MUTAG dataset, a fixed, hard-coded threshold was used instead. This inconsistency suggests possible cherry-picking and weakens reproducibility.
To ensure a fair comparison, we re-ran GLGExplainer with the elbow method enabled and all graph-level filtering disabled, ensuring that all methods were evaluated on the same test set. Under these conditions, GLGExplainer exhibited a significant performance drop compared to the original results reported in [3]. Our reproduced results are consistent with those reported in [2].
A.4 Additional experiments on classification
To ensure robust and reliable performance, we conduct additional evaluations of runtime and accuracy using three random seeds. The mean and standard deviation are reported in Table 4 and Table 5, respectively. We also report weighted precision, recall, and F1-score, averaged across the three runs, in Table 6, Table 7, and Table 8, respectively. The highest accuracy and fastest runtime among explanation methods are highlighted in bold.
Table 4: Runtime (seconds) on graph classification datasets.
Method BAShapes BBBP Mutagenicity NCI1 IMDB
GLG 272.6 $\pm$31.7 329.3 $\pm$ 46.8 744.0 $\pm$ 59.2 883.7 $\pm$ 74.5 336.6 $\pm$ 42.2
G-Trail 2,676.2 $\pm$ 124.5 5,418.9 $\pm$ 534.7 20,167.0 $\pm$ 1,047.9 25,226.1 $\pm$ 1,435.7 1,103.7 $\pm$ 67.3
$\phi_{M}$ (Ours) 50.7 $\pm$ 5.8 47.5 $\pm$ 7.4 104.7 $\pm$ 13.5 127.6 $\pm$ 15.2 67.2 $\pm$ 2.9
Table 5: Test accuracy (%) on graph classification datasets.
Method BAShapes BBBP Mutagenicity NCI1 IMDB
GNN 81.00 $\pm$ 0.41 83.33 $\pm$ 1.31 74.81 $\pm$ 1.23 75.18 $\pm$ 1.63 71.83 $\pm$ 1.43
GLG 58.83 $\pm$ 1.43 52.42 $\pm$ 1.03 57.90 $\pm$ 1.73 54.46 $\pm$ 1.85 53.83 $\pm$ 1.43
G-Trail 81.00 $\pm$ 0.41 81.13 $\pm$ 0.92 65.63 $\pm$ 0.69 66.67 $\pm$ 2.20 56.83 $\pm$ 0.62
$\phi_{M}$ (Ours) 91.00 $\pm$ 1.41 85.13 $\pm$ 2.20 76.35 $\pm$ 1.35 74.74 $\pm$ 2.12 65.67 $\pm$ 2.09
Table 6: Weighted precision (%) on graph classification datasets.
Method BAShapes BBBP Mutagenicity NCI1 IMDB
GNN 79.27 $\pm$ 1.44 82.30 $\pm$ 1.75 75.34 $\pm$ 0.98 75.19 $\pm$ 1.62 71.85 $\pm$ 1.43
GLG 58.87 $\pm$ 1.24 53.56 $\pm$ 0.62 58.48 $\pm$ 1.71 54.47 $\pm$ 1.84 53.89 $\pm$ 1.18
G-Trail 81.72 $\pm$ 0.60 79.08 $\pm$ 1.31 69.50 $\pm$ 1.32 67.06 $\pm$ 2.22 61.60 $\pm$ 1.82
$\phi_{M}$ (Ours) 92.04 $\pm$ 1.10 84.34 $\pm$ 2.44 76.41 $\pm$ 1.44 74.76 $\pm$ 2.13 70.84 $\pm$ 2.72
Table 7: Weighted recall (%) on graph classification datasets.
Method BAShapes BBBP Mutagenicity NCI1 IMDB
GNN 78.83 $\pm$ 1.70 83.33 $\pm$ 1.31 74.81 $\pm$ 1.23 75.18 $\pm$ 1.62 71.83 $\pm$ 1.43
GLG 58.83 $\pm$ 1.43 52.42 $\pm$ 1.03 57.90 $\pm$ 1.73 54.46 $\pm$ 1.85 53.83 $\pm$ 1.43
G-Trail 81.00 $\pm$ 0.41 81.13 $\pm$ 0.92 65.63 $\pm$ 0.69 66.67 $\pm$ 2.20 56.83 $\pm$ 0.62
$\phi_{M}$ (Ours) 91.00 $\pm$ 1.41 85.13 $\pm$ 2.20 76.35 $\pm$ 1.35 74.74 $\pm$ 2.12 65.67 $\pm$ 2.09
Table 8: Weighted F1-score (%) on graph classification datasets.
Method BAShapes BBBP Mutagenicity NCI1 IMDB
GNN 78.64 $\pm$ 1.80 81.86 $\pm$ 1.35 74.24 $\pm$ 1.34 75.18 $\pm$ 1.62 71.79 $\pm$ 1.38
GLG 58.73 $\pm$ 1.32 48.41 $\pm$ 1.09 58.02 $\pm$ 1.72 54.46 $\pm$ 1.84 53.74 $\pm$ 1.33
G-Trail 80.77 $\pm$ 0.38 78.55 $\pm$ 0.98 61.56 $\pm$ 0.68 66.40 $\pm$ 2.25 48.65 $\pm$ 0.91
$\phi_{M}$ (Ours) 90.89 $\pm$ 1.45 84.24 $\pm$ 2.34 76.29 $\pm$ 1.44 74.73 $\pm$ 2.12 63.44 $\pm$ 1.33
A.5 Evaluating LogicXGNN performance on additional datasets
We evaluate the performance of LogicXGNN on additional graph classification datasets to further support our claim that:
  • LogicXGNN can outperform original GNNs on datasets with well-structured, domain-specific patterns.
  • However, in cases with weaker structural regularities, there remains a performance gap between LogicXGNN and original GNNs.
To this end, we train GCNs on the HIN [3], PROTEINS [16], and AIDS [16] datasets, with dataset statistics summarized in Table 10. Among these, HIN represents a face-to-face interaction network collected in a hospital environment, PROTEINS contains large and complex compounds related to enzymes, and AIDS consists of molecular graphs labeled according to anti-HIV activity. Notably, HIN and PROTEINS exhibit weaker structural regularities, whereas AIDS contains clearer, well-defined patterns. Our results in Table 9 further reinforce our claim. Addressing the performance gap on datasets with weaker structural signals is left as an important direction for future work.
Table 9: Test accuracy (%) on additional datasets.
Method HIN PROTEINS AIDS
GNN 85.04 $\pm$ 0.58 70.85 $\pm$ 0.37 81.17 $\pm$ 0.82
G-Trail 50.57 $\pm$ 0.61 58.74 $\pm$ 1.32 87.25 $\pm$ 0.54
$\phi_{M}$ (Ours) 78.98 $\pm$ 1.29 63.97 $\pm$ 0.84 94.25 $\pm$ 0.94
Table 10: Statistics of the additional datasets.
HIN PROTEINS AIDS
#Graphs 1,760 1,113 2,000
Avg. $|\mathcal{V}|$ 11.68 39.1 15.69
Avg. $|\mathcal{E}|$ 36.06 145.6 32.39
#Node features 5 3 38
Table 11: Evaluation of LogicXGNN across different GNN architectures, measured by classification accuracy (%). $M$ and $\phi_{M}$ denote the original GNN model and its corresponding LogicXGNN-based explanation, respectively. The higher accuracy in each pair is bolded.
GCN [10] GraphSAGE [8] GIN [29] GAT [23]
Dataset $M$ $\phi_{M}$ $M$ $\phi_{M}$ $M$ $\phi_{M}$ $M$ $\phi_{M}$
BAShapes 79.00 90.50 47.00 47.00 92.00 93.50 47.00 47.00
BBBP 79.41 83.33 82.60 84.31 81.86 82.60 83.58 83.33
Mutagenicity 76.04 76.50 76.27 77.65 80.07 80.07 76.04 77.53
IMDB 71.50 65.50 73.00 66.50 73.50 66.50 71.00 65.00
NCI1 69.95 69.34 72.26 70.68 76.28 76.03 70.56 69.83
A.6 Empirical evidence for generalizability across GNN architectures
In Table 11, we present the robustness of LogicXGNN compared to the original GNN models across various architectures. LogicXGNN consistently achieves high classification accuracy and even outperforms its neural counterparts in several scenarios. These results further substantiate the strong performance of LogicXGNN in preserving predictive quality while providing interpretable explanations.
A.7 Evaluating LogicXGNN performance on challenging out-of-distribution scenarios
To further supports our claim that rule-based models can perform as well as trained GNNs, we evaluate LogicXGNN on more challenging out-of-distribution (OOD) scenarios. To this end, we implemented Murcko scaffold splitting with an 80:20 train-test ratio, repeated across three different random seeds. Table 12 reports the mean and standard deviation of test accuracies. Our evaluation uses datasets from MoleculeNet, where SMILES strings are available for scaffold computation.
Table 12: Test accuracy (%) under Murcko scaffold-based splitting on MoleculeNet datasets. Results are reported as mean $\pm$ standard deviation over three runs.
Model BACE BBBP SIDER
GNN 72.34 $\pm$ 1.18 82.52 $\pm$ 0.84 74.01 $\pm$ 0.85
LogicXGNN 74.22 $\pm$ 1.66 84.71 $\pm$ 0.93 76.11 $\pm$ 1.13
Notably, LogicXGNN outperforms the original GNN by a certain margin in this OOD setting compared to the random-split scenario. This suggests that LogicXGNN effectively captures the underlying logic governing GNN predictions. We further hypothesize that many graph tasks—particularly those involving molecules—may naturally align with logic-rule structures, which are not always well approximated by conventional GNNs. In future work, we plan to explore program synthesis approaches built upon LogicXGNN to uncover such rule-based patterns.
A.8 Rule Stability under Perturbations
We assess the robustness of extracted rules by conducting 5-fold cross-validation on multiple graph datasets, as reported in Table 13. In each fold, we independently extract rules and compared the logical forms (conjunctive clauses) against those from the baseline fold. Across all datasets, we observe a high overlap in rule structures—between 88% and 96%—demonstrating strong consistency.
This high rule matching confirms that LogicXGNN identifies stable, semantically meaningful patterns in the GNN’s decision process. Rather than overfitting to specific training samples, our method uncovers fundamental relational logic that persists across data splits, which is essential for building trust in explainable AI systems.
Table 13: 5-fold cross-validation accuracy (%) and rule overlap (%) with fold 1.
Dataset Fold 1 Fold 2 Fold 3 Fold 4 Fold 5
BAMultiShapes 88.00 (–) 86.50 (95.67) 86.25 (93.75) 87.50 (96.15) 86.00 (94.23)
Mutagenicity 76.50 (–) 75.46 (93.15) 76.15 (91.78) 75.12 (94.52) 76.61 (93.97)
BBBP 81.37 (–) 80.39 (88.10) 83.09 (92.86) 81.13 (95.24) 80.15 (90.48)
Table 14: Node classification accuracy (%) on heterogeneous graph datasets.
Method Cora PubMed
GNN 78.23 81.41
LogicXGNN 80.44 80.63
A.9 Evaluating LogicXGNN on node classification
Our approach naturally extends to node classification tasks, as it models message passing at the node level. This setting is currently unsupported by existing global explainers. We report results on two heterogeneous datasets in Table 14.
Figure 5: Histograms showing the number of instances covered by each conjunctive clause in the rule sets extracted by LogicXGNN for the BBBP and MUTAG datasets.
(a) Class 0 in BBBP (33 conjunctive clauses).
(b) Class 1 in BBBP (41 conjunctive clauses).
(c) Class 0 in MUTAG (184 conjunctive clauses).
(d) Class 1 in MUTAG (158 conjunctive clauses).
A.10 More details on knowledge extraction
We provide additional details on our knowledge extraction experiments in Section 4.2. Table 15 presents the number of extracted conjunctive clauses (patterns) for each class across different methods. Notably, LogicXGNN captures a diverse set of patterns that better reflect the intricate decision-making processes of GNNs, while baseline methods (GLGExplainer and GraphTrail) tend to underestimate this complexity.
We present complete rule sets produced by baseline approaches on Mutagenicity and BBBP in Figure 7(b). In comparison, LogicXGNN generates more diverse patterns, as quantified in Table 15 (e.g., 184 and 158 patterns for Class 0 and Class 1 in Mutagenicity, respectively). We report only the top 10 most representative conjunctive clauses in Figure 8(d), all of which are scientifically meaningful and help explain GNNs’ prediction outcomes. Selected patterns appear in the main paper (Figure 2).
Although our approach produces a large number of patterns, we observe that only a small subset of them dominate in terms of coverage, explaining the majority of graph instances. Many of the remaining patterns are highly specific and account for only a few instances, as shown in Figure 6(d).
Table 15: Number of extracted conjunctive clauses (patterns) for each class using different methods.
Dataset Class 0 Class 1
$\phi_{M}$ (Ours) G-Trail GLG $\phi_{M}$ (Ours) G-Trail GLG
Mutagenicity 184 5 1 158 1 1
BBBP 33 1 5 41 2 3
Figure 6: Complete rules produced by baseline approaches for MUTAG and BBBP.
(a) Complete rules of GraphTrail for MUTAG.
(b) Complete rules of GLGExplainer for BBBP.
Figure 7: Top 10 most representative conjunctive clauses (patterns) extracted by LogicXGNN for the MUTAG and BBBP datasets. Each pattern appears in a dedicated cell, with clauses combined via logical OR operators to form the classification rules.
(a) Rules for Class 0 in MUTAG.
(b) Rules for Class 1 in MUTAG.
(c) Rules for Class 0 in BBBP.
(d) Rules for Class 1 in BBBP.